OpenStudy (123ab456c):

Of the members of a student choir, 60% are boys. If there are 21 boys in the choir, how many members does the choir have in total? A) 33 B) 35 C) 42 D) 53

OpenStudy (anonymous):

35

OpenStudy (anonymous):

Work should look like 21*100/60=35

OpenStudy (123ab456c):

Which value for y makes this sentence true? y=1/5 A) y = 2% B) y = 20% C) y = 50% D) y = 15%

OpenStudy (anonymous):

B. y=20% remember fractions are decimals and percents too. 1/5 * 20 = 20/100 = 20%